OpenStudy (anonymous):

Evaluate the indefinite integral ∫3^2x/1+3^2x dx

OpenStudy (anonymous):

@mgeorge20 let u = 1 + 3^2x

OpenStudy (anonymous):

@mgeorge20 since the base of (u) is a constant to find du we use (rule # 12) \[\frac{ d }{ dx } [a^{u}] = (\ln(a))a ^{u}u'\]

OpenStudy (anonymous):

I should have mentioned first let's re-write it as \[\int\limits_{?}^{?}\frac{ 1 }{ 1 + 3^{2x} } * 3^{2x}dx\] to solve du of 3^2x a = 3 Let's use a different variable for 2x v = 2x du = ln(3)(3^2x)(2) we want du should equal 3^2x so that we could write an equation \[\int\limits_{?}^{?}\frac{ 1 }{ u } du\]

OpenStudy (anonymous):

@mgeorge20 the ? is not supposed to be there

OpenStudy (anonymous):

du = ln(3)(3^2x)(2) dx

OpenStudy (anonymous):

If we divide both sides by 2 ln(3) we get du/[2 ln(3)] = 3^2x dx

OpenStudy (anonymous):

@mgeorge20 \[\int\limits_{}^{} \frac{ 1 }{ u } * \frac{ du }{ 2 \ln(3) }\] 2 ln(3) is a constant so we can put in front \[\frac{ 1 }{ 2 \ln(3) } \int\limits_{}^{}\frac{ 1 }{ u }*du\]

OpenStudy (anonymous):

@mgeorge20 let me know when you're ready for the next step

OpenStudy (solomonzelman):

\(\large\color{#006699}{ \displaystyle \int\frac{a^{x} }{b+a^{x}}{~\rm d}x }\) \(\large\color{#039699}{ \displaystyle u=b+a^{x} }\) \(\large\color{#039699}{ \displaystyle a^{x}=u-b }\) \(\large\color{#039699}{ \displaystyle du/dx= \ln (a)\cdot a^{x} }\) \(\large\color{#039699}{ \displaystyle du= \ln (a)\cdot a^{x}dx }\) \(\large\color{#039699}{ \displaystyle du= \ln (a)\cdot (u-b)dx }\) \(\large\color{#039699}{ \displaystyle \frac{1}{\ln(a)}\frac{1}{ (u-b)}du= dx }\) \(\large\color{#006699}{ \displaystyle\frac{1}{\ln(a)}\cdot \int\frac{u-b}{u} \cdot\frac{1}{ (u-b)}{~\rm d}u }\) \(\large\color{#006699}{ \displaystyle\frac{1}{\ln(a)}\cdot \int\frac{1}{u} {~\rm d}u }\) \(\large\color{#006699}{ \displaystyle\frac{1}{\ln(a)}\ln(u)\color{grey}{+C} }\) \(\large\color{#006699}{ \displaystyle\frac{1}{\ln(a)}\ln(b+a^{x})\color{grey}{+C} }\)

OpenStudy (solomonzelman):

\(\large\color{#222299}{ \displaystyle \int\frac{a^{x} }{b+a^{x}}{~\rm d}x =\frac{1}{\ln(a)}\ln(b+a^{x})\color{grey}{+C} }\)

OpenStudy (anonymous):

@LeibyStrauss go ahead

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ u } = \ln \left| u \right|+ c\]

OpenStudy (solomonzelman):

(Yes.... assuming we don't get into imaginaries, the u is positie for all x anyway.

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 \ln(3) } * \ln \left| 1 + 3^{2x} \right|+ C\]

OpenStudy (anonymous):

Previous post is the final answer